OpenStudy (anonymous):

Solve by the linear combination method (with or without multiplication). 4x – 5y = –18 3x + 2y = –2

OpenStudy (anonymous):

@Compassionate

OpenStudy (smileyxl3):

2*(4x-5y)=2(-18) 5*(3x+2y)=5(-2) 8x-10y=-36 5*(3x+2y)=5(-2) 8x-10y=-36 15x+10y=-10 (15x+10y=-10)+(8x-10y=-36) 23x=-46 x=-2 Substitute that into one of the equations to get y

OpenStudy (anonymous):

Thanks, y=2